40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 1 16:01:48 2011, Jamie, Update, SUS, ETM binary whitening switching fixed 
    Reply  Fri Jul 1 18:54:34 2011, Jamie, Update, SUS, ETMY binary whitening switching confirmed to be fixed 
Message ID: 4929     Entry time: Fri Jul 1 16:01:48 2011     Reply to this: 4932
Author: Jamie 
Type: Update 
Category: SUS 
Subject: ETM binary whitening switching fixed 

I have fixed the binary whitening switching for the ETMs (ETMX and ETMY).  See below for a description of what some of the issues were.

The ETMX whitening/no-whitening response (same measurements performed in my previous post on checking vertex sus whitening switching) looks as it should.  The ETMY response seems to indicate that the switching is happening, but the measurements are very noise.  I had to up the averaging significantly to get anything sensible.  There's something else going on with ETMY.  I'll follow up on that in another post.

response

ETMX
ul : 3.28258088774 = 10.3243087313 db
ll : 3.31203559803 = 10.4018999194 db
sd : 3.27932572306 = 10.3156911129 db
lr : 3.28189942386 = 10.3225053532 db
ur : 3.31351020008 = 10.4057662366 db

ETMY
ul : 2.9802607099  =  9.4850851468 db
ll : 1.46693103911 =  3.3281939600 db
sd : 2.19178266285 =  6.8159497462 db
lr : 2.2716636118  =  7.1268804285 db
ur : 3.42348315519 = 10.6893639064 db

End rack cable diagrams inconsistent with binary channel mapping

One of the big problems was that the most up-to-date end rack cable diagrams (that I can find) are inconsistent with the actual binary mapping. The diagram says that:

  • BO adapter chassis output A (ch 1-16)   --> CAB_1X4_26 --> cross-connect 1X4-B7 (carrying QPD whitening switching signals)
  • BO adapter chassis output B (ch 17-32) --> CAB_1X4_27 --> cross-connect 1X4-A6 (carrying OSEM whitening switching signals)

In fact, the binary outputs are switched, such that output A carries the OSEM signals, and output B carries the QPD whitening signals.

I SWITCHED THE CABLES AT THE BINARY OUTPUT ADAPTER CHASSIS so that:

  • BO adapter chassis output A (ch 1-16)   --> CAB_1X4_27 --> cross-connect 1X4-A6 (carrying OSEM whitening switching signals)
  • BO adapter chassis output B (ch 17-32) --> CAB_1X4_26 --> cross-connect 1X4-B7 (carrying QPD whitening switching signals)

The rest of the wiring remains the same.

I made the same transformation for ETMY as well.

ELOG V3.1.3-